Material, Labor, and Variable Overhead Variances
The following summarized manufacturing data relate to Thomas Corporation’s April operations, during which 2,000 finished units of product were produced. Normal monthly capacity is 1,100 direct labor hours.

Standard Units Costs Total Actual Costs
Direct material
Standard (2 lb. @ $9.00/lb.) $18
Actual (4,200 lb. @ $10.20/lb.) $42,840
Direct labor
Standard (0.5 hr. @ $24/hr.) $12
Actual (950 hrs. @ $23.40/hr.) 22,230
Variable overhead
Standard (0.5 hr. @ $6/hr.) $3
Actual - 6,450
Total $33 $71,520

Determine the following variances:

Do not use negative signs with any of your answers. Next to each variance answer, select either "F" for Favorable or "U" for Unfavorable.

Materials Variances
Actual cost: Answer
Split cost: Answer
Standard cost: Answer
Materials price Answer AnswerFU
Materials efficiency Answer AnswerFU
Labor Variances
Actual cost: Answer
Split cost: Answer
Standard cost: Answer
Labor rate Answer AnswerFU
Labor efficiency Answer AnswerFU
Variable Overhead Variances
Actual cost: Answer
Split cost: Answer
Standard cost: Answer
Variable overhead spending Answer AnswerFU
Variable overhead efficiency Answer AnswerFU
Answer #1

Material Variance

Actual Cost $ 10.20
Split Cost
Standard Cost $ 9
Material Price( $ 9 - $ 10.20 = $1.20 × 4,200) $5040 U
Materials Efficency ( Actual usage - Standard usage )(4200 - 4000) $ 200 U

Labour Variance

Actual Cost $ 23.40
Split Cost
Standard Cost $24
Labour Rate ( 24 - 23.40 = .60 * 950 = $ 570) $570 F
Labour Efficiency ( .50*2000 = $1000 - 950 = 50 * 24 = $ 1200) $1,200 F

Labour rate is favourable due to 60 cent savings

Variable Overhead Variance

Actual Cost Not Given
Split Cost
Standard Cost $ 6
Variable Overhead Spending ($950 * 6 = $ 5,700 , $6450 - 5700 = $ 750 $ 750 F
Variable Overhead Efficiency ( $1000 - 950) * 6 = $ 300 F $ 300 F